A coin tossed and a die is rolled. The outcome of the coin is recorded 1 when it show a head and 0 when it shows a tail. The random variable R gives the sum of the outcomes of the coin and the die. Compute the average value of the random variables.
1
Expert's answer
2022-02-18T06:10:56-0500

The sample space "S" is given below,

"S=\begin{Bmatrix}
 2 & 1 \\
 3 & 2\\
4&3\\
5&4\\
6&5\\
7&6
\end{Bmatrix}"

The probability distribution is,

"x"  1 2 3 4 5 6 7

"p(x)" "{1\over 12}" "{2\over12}" "2\over12" "2\over12" "2\over12" "2\over12" "1\over12"

The expected value is given as,

"E(x)=\sum xp(x)=(1\times {1\over12})+(2\times {2\over12})+(3\times {2\over12})+(4\times {2\over12})+(5\times {2\over12})+(6\times {2\over12})+(7\times {1\over12})=4" Therefore, the average of the random variable is 4.